OpenStudy (anonymous):

the dimensions of a rectangle are such that its length is 5 in. more than its width. If the length were doubled and if the width were decreased by 2 in., the area would be increased by 162 in squared. What are the length and width of the rectangle?

OpenStudy (karatechopper):

New dimensions: width = x+2 length = 2(x+5) Area is 2(x+5)(x+2) = 2(x^2+7x+10) = 2x^2+14x+20

OpenStudy (karatechopper):

Original dimensions: Let the width be "x". Then the length is "x+5" Area is x(x+5) = x^2+5x.